Ball A with mass of 100 g moving in a constant velocity of 200 cm/s,hits ball B with a mass of 400 g which is at rest. As a result of the collision, Ball A bounced back at 120 cm/s. Determine the final velocity of ball B


1
Expert's answer
2021-11-07T19:27:17-0500

Momentum conservation:


"m_Av_A+0=m_Au_A+m_Bu_B,\\\space\\
u_B=\frac{m_Av_A-m_Au_A}{m_B},\\\space\\
u_B=\frac{0.1·2-0.1·(-1.2)}{0.4}=0.8\text{ m/s}"